OpenStudy (anonymous):

In Los Angeles, the tax on a property assessed at $910,000 is $20,020. If tax rates are proportional in this city, how much would the tax be on a property assessed at $950,000?

OpenStudy (anonymous):

To answer this question we must use Solve for x. So, if a $910,000 property had a tax of 20,020, \[\frac{ x $}{ 950,000$ } = \frac{ 20,020 $}{ 910,000$ }\] \[x = \frac{ 20,020 $}{ 910,000$ } \times \frac{ 950,000 }{ 1 }\] \[x = \frac{ $19,019,000.00 }{ 910,000}\] \[X =\frac{ 20,900 }{ 1}\]

OpenStudy (anonymous):

So, the tax would be $20,900.